OpenStudy (anonymous):

A scientist counts 35 bacteria present in a culture and finds that the number of bacteria triples each hour. The function y = 35 • 3^x models the number of bacteria after x hours. Graph the function. (You do not need to submit the graph). Use the graph to estimate when there will be about 550 bacteria in the culture.

Directrix (directrix):

Isn't this also from night before last? That is, hasn't it been answered?

OpenStudy (anonymous):

i answered it but someone got me confused with log and In

Directrix (directrix):

Was it this thread where you became confused? http://openstudy.com/users/lilyswan#/updates/51671858e4b054d007825f28

OpenStudy (anonymous):

yep the same

Directrix (directrix):

Did you begin with 500 = 35 * 3^x to solve for x? (I know you are supposed to read it from the graph but if the question has no answer options, reading values from a hand drawn graph of an exponential function is dicey.

OpenStudy (anonymous):

sorry for confusing you

OpenStudy (anonymous):

now i started it with 550 = 35*3^x

OpenStudy (anonymous):

its alright peter i just want to clear my confusion its not because of you to be true log and in really confuses me at time

OpenStudy (anonymous):

and this question is the tricky one

Directrix (directrix):

550 = 35*3^x (1/35) * 550 = (1/35) * 3^x Do you agree with that. I'm dividing by 35 as did the others.

OpenStudy (anonymous):

ok

Directrix (directrix):

15.714286 (approx) = 3^x Agree?

Directrix (directrix):

(1/35) * 550 = 15.714286 (approx)

Directrix (directrix):

@LilySwan

OpenStudy (anonymous):

yea i got

Directrix (directrix):

log (15.714286) = log (3^x) --> These are base 10 logs.

Directrix (directrix):

Agree?

OpenStudy (anonymous):

ok

Directrix (directrix):

log (15.714286) = log (3^x) log (15.714286) = x * (log 3)

OpenStudy (anonymous):

1.19=x^0.477

Directrix (directrix):

Why log (3^x) = x * (log 3)

Directrix (directrix):

1.19=x^0.477 --> Try not to jump ahead. Be a step-sister. Show all the steps.

OpenStudy (anonymous):

ok

Directrix (directrix):

log (15.714286) = x * (log 3) 1.19629465304 = x * (log 3)

Directrix (directrix):

1.19629465304 = x * 0.47712125472

OpenStudy (anonymous):

ok

Directrix (directrix):

1.19629465304 divided by .47712125472 = x

Directrix (directrix):

2.50731788032 = x

OpenStudy (anonymous):

ok so we will use log

OpenStudy (anonymous):

thanks trix for clearing my doubts

Directrix (directrix):

x is roughly 2.5 hours required to get 550 bacteria.

OpenStudy (anonymous):

thanks

Directrix (directrix):

Okay, then. Ha sido un placer

OpenStudy (anonymous):

ill be back i have to take my medicine in 10 minutes